# Topology and Diffeology via Metric-like Functions

### v1

@@ 2. Topology defined by $d_{r}$
To prove the inclusion, let $g\in\mathcal{U}_{r,\varepsilon}(f)$ and take any $x\in K$. Since $x$ belongs to one of the preimages $f^{-1}(B_{\varepsilon_{x_{i}}/2}{(f(x_{i}))})$, we have
@@
To prove the inclusion, let $g\in\bigcap_{i=1}^{n}W_{i}$ and take $x\in{r}D$. Since $x$ is contained in some $B_{r_{x_{i}}}{(x_{i})}$, we have $g(x)\in B_{\varepsilon/2}{(f(x_{i}))}$. Moreover, since $B_{r_{x_{i}}}{(x_{i})}\subset f^{-1}(B_{\varepsilon/2}{(f(x_{i}))})$, we obtain $f(x)\in B_{\varepsilon/2}{(f(x_{i}))}$. Thus,
@@ 3. Tiling spaces and $d_{r}$
Take any $r>0$, $T_{1},T_{2}\in\Omega_{T}$, and $\varepsilon>0$ such that $d_{r}(T_{1},T_{2})<\varepsilon$. Then, there exists $u\in\mathbb{R}^{d}$ such that $|u|<\varepsilon$, and $(T_{1}+u)\cap{r}D=T_{2}\cap{r}D$. Since the set
@@
Finally, we prove the weak triangle inequality. Take any $T_{1},T_{2},T_{3}\in\Omega_{T}$, and let $r_{1},r_{2},r_{3}>0$ satisfy $d_{r_{12}}(T_{1},T_{2})<r_{3},\quad d_{r_{123}}(T_{2},T_{3})<r_{2}$. Also, take any $d_{1},d_{2}>0$ such that
@@
The family $\{d_{r}\}_{r>0}$ not only induces a topology but also a quasi-uniformity. Moreover, an even weaker version of the triangle inequality suffices to define the quasi-uniformity. In particular, this weaker triangle inequality is satisfied in the space of graph embeddings $\Phi{(\mathbb{R}^{N})}$, as defined in [6].
@@
We now introduce an even weaker version of the weak triangle inequality.
@@
A set $(X,\{d_{r}\}_{r>0})$ is said to satisfy the weaker $\{d_{r}\}$ axioms if it satisfies the axioms in Definition 2.1, except that the weak triangle inequality is replaced by the weaker triangle inequality.
@@
Let $(X,\{d_{r}\}_{r>0})$ satisfy the weaker $\{d_{r}\}$ axioms in Definition 4.4. We define
@@
for each $r>0$. If $\mathcal{P}_{r}(G,H)=\emptyset$, we set $d_{r}(G,H)=\infty$. This family $\{d_{r}\}_{r>0}$ satisfies the weaker $\{d_{r}\}$ axioms in Definition 4.4. The topology induced by $\{d_{r}\}_{r>0}$ coincides with the $C^{0}$-topology defined in [6], Definition 2.5.
@@
Take any $r>0$, $G,H\in\Psi_{d}(\mathbb{R}^{N})$, and $\varepsilon>0$ such that $d_{r}(G,H)<\varepsilon$. Then, there exists a map $\varphi\in\mathcal{P}_{r}(G,H)$ with $\varphi\colon U\to\phi(U)$ satisfying $|p-\varphi(p)|<\varepsilon$ for all $p\in s(\varphi,r)$. Since $U$ is open in $H$, the set
@@
Take any $G_{1},G_{2},G_{3}\in\Phi(\mathbb{R}^{N})$ and any $r_{1},r_{2},r_{3}>0$. Suppose that
@@
For this composition $\varphi_{1}\circ\varphi_{2}$, we have $\varphi_{1}\circ\varphi_{2}\in\mathcal{P}_{r_{1}}(G_{1},G_{3})$. Indeed, take any $p_{1}\in G_{1}\cap{r_{1}}D\subset G_{1}\cap{r_{123}}D\subset\varphi_{1}(U_{1})$. There exists $p_{2}\in U_{1}$ such that $p_{1}=\varphi_{1}(p_{2})$. Since $|\varphi_{1}(p_{2})-p_{2}|<d_{3}<r_{3}$ holds, we have $p_{2}\in G_{2}\cap{r_{13}}D\subset\varphi_{2}(U)$. Therefore, we obtain $G_{1}\cap{r_{1}}D\subset\varphi_{1}\circ\varphi_{2}(U)$.
@@
Let $(X,\{d_{r}\}_{r>0})$ satisfy the weaker $\{d_{r}\}$ axioms in Definition 4.4. A parametrization $p\colon U_{p}\to X$ is smooth at $t_{0}\in U_{p}$ if there exists $\delta>0$ satisfying the following conditions:
@@
Since the covering and locality axioms follow directly from the definitions, we proceed to prove the smooth compatibility axiom. Let $p\colon U_{p}\to X$ be a plot of $X$, and let $f\colon V\to U_{p}$ be a smooth map from a Euclidean open set. Take any $t_{0}\in V$. Since $p$ is smooth at $f(t_{0})$, there exists $\delta>0$ such that the following conditions hold:
@@
First, we prove that $F_{T_{1}}$ is a plot of the diffeology induced by $\{d_{r}\}_{r>0}$ for each $T_{1}\in\Omega_{T}$. Let $\mathscr{P}$ be the set of prototiles of $T$, and let $T_{1}\in\Omega_{T}$ a tiling. Take any $t_{0}\in\mathbb{R}^{d}$, and define $\delta=\min_{\tau\in\mathscr{P}}(\mathop{\mathrm{diam}}{\tau})$. Then, for each $t\in B_{\delta}{(t_{0})}$, we have
@@
On the other hand, since $t\in\overline{A}_{r^{\prime}}$, there exists a sequence $\{s_{n}\}_{n\in\mathbb{N}}$ converging to $t$ such that
@@
Therefore, for any $\varepsilon^{\prime}>0$, there exist $\delta^{\prime}>0$ such that for each $t\in B_{\delta^{\prime}}{(t_{0})}$,
@@ 6. Relating topologies and diffeologies
Let $(X,\{d_{r}\}_{r>0})$ satisfy the weaker $\{d_{r}\}$ axioms in Definition 4.4. If a map $p\colon U_{p}\to X$ is a plot of $X$ as defined in Proposition 5.6, then $p$ is also continuous.
@@
Let $(X,\{d_{r}\}_{r>0})$ satisfy the weaker $\{d_{r}\}$ axioms in Definition 4.4. The $D$-topology of the diffeology on $X$ induced by the family $\{d_{r}\}_{r>0}$ (see Proposition 5.6) is finer than the topology on $X$ induced by the same family (see Proposition 4.6).

### v2

@@ 1. Introduction
We also show that a family $\{d_{r}\}_{r>0}$ gives rise to a natural pseudo-metric that is not necessarily symmetric on the space. In particular, in the case of submanifolds, this provides a simpler way to recover the result of [3], Theorem 4.4.
@@ 2. Topology defined by $d_{r}$
To prove the inclusion, let $g\in\mathcal{U}_{r,\varepsilon}(f)$ and $x\in K$ be arbitrary. Since $x$ belongs to one of the preimages $f^{-1}(B_{\varepsilon_{x_{i}}/2}{(f(x_{i}))})$, we have
@@
To prove the inclusion, let $g\in\bigcap_{i=1}^{n}W_{i}$ and $x\in{r}D$. Since $x$ is contained in some $B_{r_{x_{i}}}{(x_{i})}$, we have $g(x)\in B_{\varepsilon/2}{(f(x_{i}))}$. Moreover, since $B_{r_{x_{i}}}{(x_{i})}\subset f^{-1}(B_{\varepsilon/2}{(f(x_{i}))})$, we obtain $f(x)\in B_{\varepsilon/2}{(f(x_{i}))}$. Thus,
@@ 3. Tiling spaces and $d_{r}$
Let $r>0$, $T_{1},T_{2}\in\Omega_{T}$, and $\varepsilon>0$ be arbitrary satisfying $d_{r}(T_{1},T_{2})<\varepsilon$. Then, there exists $u\in\mathbb{R}^{d}$ such that $|u|<\varepsilon$, and $(T_{1}+u)\cap{r}D=T_{2}\cap{r}D$. Since the set
@@
Finally, we prove the weak triangle inequality. Let $T_{1},T_{2},T_{3}\in\Omega_{T}$, and $r_{1},r_{2},r_{3}>0$ be arbitrary satisfying $d_{r_{12}}(T_{1},T_{2})<r_{3}$, $d_{r_{123}}(T_{2},T_{3})<r_{2}$. Also, let $d_{1},d_{2}>0$ be arbitrary satisfying
@@ 4. Uniformity defined by $d_{r}$
The family $\{d_{r}\}_{r>0}$ induces not only a topology but also a quasi-uniformity. Moreover, an even weaker version of the triangle inequality suffices to define the quasi-uniformity. In particular, this weaker triangle inequality is satisfied in the space of graph embeddings $\Phi{(\mathbb{R}^{N})}$, as defined in [7].
@@
We next introduce a still weaker form of the triangle inequality, which generalizes the weak triangle inequality of Definition 2.1.
@@
A set $(X,\{d_{r}\}_{r>0})$ is said to satisfy the weaker $\{d_{r}\}$-axioms if it satisfies the axioms in Definition 2.1, except that the weak triangle inequality is replaced by the weaker triangle inequality.
@@
Let $(X,\{d_{r}\}_{r>0})$ satisfy the weaker $\{d_{r}\}$-axioms in Definition 4.4. We define
@@
for each $r>0$. If $\mathcal{P}_{r}(G,H)=\emptyset$, we set $d_{r}(G,H)=\infty$. This family $\{d_{r}\}_{r>0}$ satisfies the weaker $\{d_{r}\}$-axioms in Definition 4.4. The topology induced by $\{d_{r}\}_{r>0}$ coincides with the $C^{0}$-topology defined in [7], Definition 2.5.
@@
Let $r>0$, $G,H\in\Psi_{d}(\mathbb{R}^{N})$, and $\varepsilon>0$ be arbitrary satisfying $d_{r}(G,H)<\varepsilon$. Then, there exists a map $\varphi\in\mathcal{P}_{r}(G,H)$ with $\varphi\colon U\to\phi(U)$ satisfying $|p-\varphi(p)|<\varepsilon$ for all $p\in s(\varphi,r)$. Since $U$ is open in $H$, the set
@@
Let $G_{1},G_{2},G_{3}\in\Phi(\mathbb{R}^{N})$, and $r_{1},r_{2},r_{3}>0$ be arbitrary. Suppose that
@@
For this composition $\varphi_{1}\circ\varphi_{2}$, we have $\varphi_{1}\circ\varphi_{2}\in\mathcal{P}_{r_{1}}(G_{1},G_{3})$. Indeed, let $p_{1}\in G_{1}\cap{r_{1}}D\subset G_{1}\cap{r_{123}}D\subset\varphi_{1}(U_{1})$ be arbitrary. There exists $p_{2}\in U_{1}$ such that $p_{1}=\varphi_{1}(p_{2})$. Since $|\varphi_{1}(p_{2})-p_{2}|<d_{3}<r_{3}$ holds, we have $p_{2}\in G_{2}\cap{r_{13}}D\subset\varphi_{2}(U)$. Therefore, we obtain $G_{1}\cap{r_{1}}D\subset\varphi_{1}\circ\varphi_{2}(U)$.
@@
The family $\{d_{r}\}_{r>0}$ induces a pseudo-metric that is not necessarily symmetric.
Definition 4.9.
Let $X$ be a set. A quasi-pseudo-metric on $X$ is a function $d\colon X\times X\to\mathbb{R}_{\geq 0}$ satisfying the following:
Self-distance Axiom:
$d(x,x)=0$ for all $x\in X$.
Triangle inequality:
$d(x,z)\leq d(x,y)+d(y,z)$ for all $x,y,z\in X$.
If the following condition is also satisfied, then we call $d$ a quasi-metric:
Non-degeneracy:
$d(x,y)=0$ implies $x=y$.
Given a quasi-pseudo-metric $d$ on a set $X$, we define
$V_{\varepsilon}=\{(x,y)\in X\times X\mid d(x,y)<\varepsilon\}$
for each $\varepsilon>0$. A subset $U\subset X\times X$ is called an entourage if there exists $\varepsilon>0$ such that $V_{\varepsilon}\subset U$. Then, the collection $\mathcal{U}$ of all such entourages forms a quasi-uniformity on $X$.
Theorem 4.10.
Let $(X,\{d_{r}\}_{r>0})$ satisfy the weaker $\{d_{r}\}$-axioms in Definition 4.4. Then there exists a quasi-pseudo-metric $d$ on $X$ that induces the quasi-uniformity defined in Proposition 4.6. Moreover, if the family $\{d_{r}\}_{r>0}$ is non-degenerate, i.e.,
$d_{r}(x,y)=0\text{ for all }r>0\implies x=y,$
then $d$ can be chosen to be a quasi-metric on $X$.
Proof.
The following argument is based on Proposition 2 in Chapter IX, Section 1.4 of [5] by Bourbaki.
We define $W_{n}=U_{3^{2n},3^{-n}}=\{(x,y)\in X\times X\mid d_{3^{2n}}(x,y)<3^{-n}\}$ for each $n\in\mathbb{Z}_{\geq 0}$. We now show that the following inclusion holds for each $n\in\mathbb{Z}_{\geq 0}$:
$W_{n+1}\circ W_{n+1}\circ W_{n+1}\subset W_{n}.$
Let $(x,w)\in W_{n+1}\circ W_{n+1}\circ W_{n+1}$ be arbitrary. Then, there exist $y,z\in X$ such that $(x,y)$, $(y,z)$, and $(z,w)$ are all in $W_{n+1}$. By applying the weaker triangle inequality repeatedly, we obtain
$d_{3^{2n}}(x,w)\leq d_{3^{2n+1}}(x,y)+d_{3^{2n+1}}(y,w)\leq d_{3^{2n+2}}(x,y)+d_{3^{2n+1}}(y,w)$
$\leq d_{3^{2n+2}}(x,y)+d_{3^{2n+2}}(y,z)+d_{3^{2n+2}}(z,w)\leq 3^{-n-1}+3^{-n-1}+3^{-n-1}=3^{-n}.$
Using these $\{W_{n}\}_{n\in\mathbb{Z}_{\geq 0}}$, we define a function $g\colon X\times X\to\mathbb{R}_{\geq 0}$ by
$g(x,y)=\begin{cases}0&\text{if }(x,y)\in W_{n}\text{ for all }n\geq 1,\\ 2^{-k}&\text{if }(x,y)\in W_{n}\text{ for }1\leq n\leq k\text{ and }(x,y)\notin W_{k+1},\\ 1&\text{if }(x,y)\notin W_{1}.\end{cases}$
for $(x,y)\in X\times X$. Also, we define a function $d\colon X\times X\to\mathbb{R}_{\geq 0}$ by
$d(x,y)=\inf\sum_{i=0}^{p-1}g(z_{i},z_{i+1}),$
where the infimum is taken over all finite sequences $\{z_{i}\}_{0\leq i\leq p}$ such that $z_{0}=x$ and $z_{p}=y$. We shall show that $d$ is a quasi-pseudo-metric on $X$ which satisfies the following inequalities:
$\dfrac{1}{2}\,g(x,y)\leq d(x,y)\leq g(x,y).$
It is straightforward to verify that $d$ is a quasi-pseudo-metric on $X$. Since the inequality $d(x,y)\leq g(x,y)$ follows directly from the definition, we only need to prove the inequality $\frac{1}{2}g(x,y)\leq d(x,y)$. To this end, we prove the following by induction on $p\in\mathbb{Z}_{>0}$:
For any $x,y\in X$ and any finite sequence $\{z_{i}\}_{0\leq i\leq p}$ in $X$ with $z_{0}=x$ and $z_{p}=y$, we have
$\dfrac{1}{2}\,g(x,y)\leq\sum_{i=0}^{p-1}g(z_{i},z_{i+1}).$
Suppose that the statement holds for every $p^{\prime}<p$. We now prove the case $p$. We define $a=\sum_{i=0}^{p-1}g(z_{i},z_{i+1})$. If $a\geq 1/2$, then the desired inequality holds trivially. If $a<1/2$, let $h$ be the largest integer $q\in\mathbb{Z}_{\geq 1}$ such that
$\sum_{i=1}^{q-1}g(z_{i},z_{i+1})\leq\dfrac{a}{2}.$
By the definition of $h$, we have
$\sum_{i=0}^{h-1}g(z_{i},z_{i+1})\leq\dfrac{a}{2}\quad\text{and}\quad\sum_{i=0}^{h}g(z_{i},z_{i+1})>\dfrac{a}{2}.$
It follows from the second inequality that $\sum_{i=h+1}^{p-1}g(z_{i},z_{i+1})\leq a/2$. By the inductive hypothesis, we have $g(x,z_{h})\leq a$ and $g(z_{h+1},y)\leq a$. It is also clear that $g(z_{h},z_{h+1})\leq a$. Let $k\in\mathbb{Z}_{\geq 1}$ be the smallest integer such that $2^{-k}\leq a$. Since $g(x,z_{h})$, $g(z_{h},z_{h+1})$, and $g(z_{h+1},y)$ are each of the form $2^{-n}$ and less than or equal to $a$, it follows that $(x,z_{h}),\ (z_{h},z_{h+1})$, and $(z_{h+1},y)$ all belong to $W_{k}$. Therefore, we obtain $(x,y)\in W_{k}\circ W_{k}\circ W_{k}\subset W_{k-1}$, which implies that $g(x,y)\leq 2^{1-k}\leq 2a$. This completes the proof of $\frac{1}{2}\,g(x,y)\leq d(x,y)$.
Using these inequalities, we now show that the quasi-uniformity $\mathcal{U}$ induced by the family $\{d_{r}\}_{r>0}$ coincides with the quasi-uniformity $\mathcal{U}^{\prime}$ induced by $d$. Let $U\in\mathcal{U}$ be an arbitrary entourage. Then there exist $r>0$ and $\varepsilon>0$ such that $U_{r,\varepsilon}\subset U$. For these $r$ and $\varepsilon$, choose $k\in\mathbb{Z}_{\geq 0}$ such that $3^{2k}>r$, $3^{-k}<\varepsilon$. Since $g(x,y)\leq 2d(x,y)$ for any $x,y\in X$, we have
$V_{2^{-k-1}}\subset W_{k}\subset U_{r,\varepsilon}\subset U.$
Indeed, for any $(x,y)\in V_{2^{-k-1}}$, we have $g(x,y)\leq 2d(x,y)<2^{-k}$, which implies $(x,y)\in W_{k}$. Therefore, $U$ is also in $\mathcal{U}^{\prime}$.
Conversely, let $U\in\mathcal{U}^{\prime}$ be an arbitrary entourage. Then there exists $\varepsilon>0$ such that $V_{\varepsilon}\subset U$. Choose $k\in\mathbb{Z}_{\geq 0}$ such that $2^{-k}<\varepsilon$. Since $d(x,y)\leq g(x,y)$ for any $x,y\in X$, we have
$U_{3^{2k},3^{-k}}=W_{k}\subset V_{\varepsilon}.$
Indeed, for any $(x,y)\in W_{k}$, we have $d(x,y)\leq g(x,y)\leq 2^{-k}<\varepsilon$. Therefore, $U$ is also in $\mathcal{U}$.
Finally, we prove that $d(x,y)=0$ implies $x=y$ under the assumption that the family $\{d_{r}\}_{r>0}$ is non-degenerate. Let $x,y\in X$ be arbitrary with $d(x,y)=0$, and let $r>0$ and $\varepsilon>0$ be arbitrary. Choose $k\in\mathbb{Z}_{\geq 1}$ such that $W_{k}\subset U_{r,\varepsilon}$. Since $g(x,y)\leq 2d(x,y)=0$, we obtain $(x,y)\in W_{n}$ for all $n\in\mathbb{Z}_{\geq 1}$, and in particular $(x,y)\in W_{k}$. Hence, $(x,y)\in U_{r,\varepsilon}$, it follows that $d_{r}(x,y)<\varepsilon$. Since $r$ and $\varepsilon$ were arbitrary, this means $d_{r}(x,y)=0$ for all $r>0$. By the non-degeneracy of $\{d_{r}\}_{r>0}$, we conclude that $x=y$. ∎
Remark 4.11.
If each $d_{r}$ is symmetric, then the family $\{d_{r}\}_{r>0}$ induces a genuine uniformity rather than just a quasi-uniformity. In this case, Proposition 2 in Chapter IX, Section 1.4 of [5] applies directly and ensures the existence of a pseudo-metric $d$ that induces the original uniformity. Moreover, if each $d_{r}$ is symmetric, then the topology induced by the family $\{d_{r}\}_{r>0}$ satisfies the $T_{1}$ separation axiom. In fact, it is even regular ([3], Lemma 4.3). As a consequence, the pseudo-metric $d$ constructed in the proof must actually be a metric.
Example 4.12.
The family $\{d_{r}\}_{r>0}$ on the space of submanifolds $\Psi_{d}(\mathbb{R}^{N})$ in Example 2.7 is symmetric. Therefore, by Theorem 4.10 and Remark 4.11, there exists a metric on $\Psi_{d}(\mathbb{R}^{N})$ which induces both the uniformity and the topology on $\Psi_{d}(\mathbb{R}^{N})$. This provides a simpler way to recover the result of [3], Theorem 4.4.
@@ 5. Diffeology defined by $d_{r}$
Let $(X,\{d_{r}\}_{r>0})$ satisfy the weaker $\{d_{r}\}$-axioms in Definition 4.4. A parametrization $p\colon U_{p}\to X$ is smooth at $t_{0}\in U_{p}$ if there exists $\delta>0$ satisfying the following conditions:
@@
Since the covering and locality axioms follow directly from the definitions, we proceed to prove the smooth compatibility axiom. Let $p\colon U_{p}\to X$ be a plot of $X$, and let $f\colon V\to U_{p}$ be a smooth map from a Euclidean open set. Let $t_{0}\in V$ be arbitrary. Since $p$ is smooth at $f(t_{0})$, there exists $\delta>0$ such that the following conditions hold:
@@
First, we prove that $F_{T_{1}}$ is a plot of the diffeology induced by $\{d_{r}\}_{r>0}$ for each $T_{1}\in\Omega_{T}$. Let $\mathscr{P}$ be the set of prototiles of $T$, and let $T_{1}\in\Omega_{T}$ a tiling. Let $t_{0}\in\mathbb{R}^{d}$ be arbitrary, and define $\delta=\min_{\tau\in\mathscr{P}}(\mathop{\mathrm{diam}}{\tau})$. Then, for each $t\in B_{\delta}{(t_{0})}$, we have
@@
On the other hand, since $t\in\overline{A}_{r^{\prime}}$, there exists a sequence $\{s_{n}\}_{n\in\mathbb{Z}_{\geq 0}}$ converging to $t$ such that
@@
Therefore, for any $\varepsilon^{\prime}>0$, there exists $\delta^{\prime}>0$ such that for each $t\in B_{\delta^{\prime}}{(t_{0})}$,
@@ 6. Relating topologies and diffeologies
Let $(X,\{d_{r}\}_{r>0})$ satisfy the weaker $\{d_{r}\}$-axioms in Definition 4.4. If a map $p\colon U_{p}\to X$ is a plot of $X$ as defined in Proposition 5.6, then $p$ is also continuous.
@@
Let $(X,\{d_{r}\}_{r>0})$ satisfy the weaker $\{d_{r}\}$-axioms in Definition 4.4. The $D$-topology of the diffeology on $X$ induced by the family $\{d_{r}\}_{r>0}$ (see Proposition 5.6) is finer than the topology on $X$ induced by the same family (see Proposition 4.6).